slove the equation 80=3y+2y+4+1
Found 2 solutions by checkley71, kietra:
Answer by checkley71(8403)

80=3y+2y+4+1 or 80=5y+5 or 75=5y or y=15
proof 80=3*15+2*15+4+1 or 80=45+30+5 or 80=80

Answer by kietra(57)

80 = 3y + 2y + 4 + 1
First combine what is the same. We can combine the numbers on the left as well as the 3y and 2y.
80 = 3y + 2y + 4 + 1
80 = 3y + 2y + 5
80 = 5y + 5
Now put the numbers on the same side by subtracting 5 from both sides
80 - 5 = 5y + 5 - 5
75 = 5y
Divide both sides by 5 to get y by itself.
75 = 5y
75/5 = 5/5y
13 = 1y
13 = y